40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13784     Entry time: Tue Apr 24 11:31:59 2018     Reply to this: 13786
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: Proposed changes to EX fiber coupling 

Motivation: I want to make another measurement of the out-of-loop ALS beat noise, with improved MM into both the PSL and EX fibers and also better polarization control. For this, I want to make a few changes at the EX table. 

  1. Replace existing fiber collimator with one of the recently acquired F220-APC-1064 collimators.
    • This gives an output mode of diameter 2.4mm with a beam divergence angle of 0.032 degrees (all numbers theoretical - I will measure these eventually but we need a beam path of ~5m length in order to get a good measurement of this collimated beam).
    • I believe it will be easier to achieve good mode matching into this mode rather than with the existing collimator. 
    • Unfortunately, the mount is still going to be K6X and not K6XS. 
  2. Improve mode-matching into fiber.
    • I used my measurement of the Innolight NPRO mode from 2016, a list of available lenses, and some measured distances to calculate a solution that gives theoretical 100% overlap with the collimator mode, that has beam diameter 2.4mm, located 80cm from the NPRO shutter head location (see Attachment #1).
    • The required movement of components is schematically illustrated in Attachment #2.
    • One of the required lens positions is close to the bracket holding the enclosure to the table, but I think the solution is still workable (the table is pretty crowded so I didn't bother too much with trying to find alternative solutions as all of them are likely to require optics placed close to existing ones and I'd like to avoid messing with the main green beam paths.
    • I will attempt to implement this and see how much mode matching we actually end up getting.
  3. Install a PBS + HWP combo in the EX fiber coupling path.
    • This is for better polarization control.
    • Also gives us more control over how much light is coupled into the fiber in a better way than with the ND filters in current path.
  4. Clean EX fiber tip.
  5. Dump a leakage IR beam from the harmonic separator post doubling oven, which is currently just hitting the enclosure. It looks pretty low power but I didn't measure it.
  6. Re-install EX power monitoring PD.

Barring objections, I will start working on these changes later today.

Attachment 1: EX_fiber_MM.pdf  48 kB  Uploaded Tue Apr 24 12:49:31 2018  | Hide | Hide all
EX_fiber_MM.pdf
Attachment 2: EX_fiber_changes.png  7.338 MB  Uploaded Tue Apr 24 12:49:42 2018  | Hide | Hide all
EX_fiber_changes.png
ELOG V3.1.3-